FEL Monthly Highlights
June 2001

We were pleased to provide the local arrangements for the JTO sponsored workshop on “High Power FELs for Naval Applications” that was held June 5-6  in Newport News.  Over 120 participated.  The conference summary will be prepared by the workshop chair, Alan Todd from AES and distributed to attendees and interested parties within JTO and the Navy.  The bottom line: there appears to be a credible development path for 100 kW and above based on the FEL concept demonstrated with the 1 and 10 kW Demos at Jefferson Lab.  Such a program will benefit from DOD experience with other high energy laser systems and the DOE accelerator programs which are continuing to develop relevant electron sources, acceleration systems and RF power systems.

The first user operations for the Summer run provided FEL light to the nanotube collaboration from CWM, NASA-Langley and Penn State.

This month we completed the first half of the parametric study on laser-materials damage (i.e.,"lethality") funded by the Joint Technology Office.  See the "Operations" section for details.

On June 26 we hosted a review of the FEL program for SURA's Maritime Technical Advisory Committee (MTAC), chaired by VADM (ret.) Al Baciocco.

We provided updated report material to John Albertine who is editing the report from the June 7 DoD FEL Technology Area Working Group meeting.  Our comments included our best current estimates for an FEL development plan for scaling FELs to 100 kW and then to the MW level.

Several papers related to the FEL design and performance were presented the 2001 Particle Accelerator Conference in Chicago this month.  Dave Douglas presented an overview on the FEL Upgrade design (see next section for the reference); Steve Benson presented a paper on FEL harmonic generation; Lia Merminga presented a paper on high average current effects in energy recovered linacs; and Charlie Sinclair presented a paper on our ion implantation procedures for high voltage electrodes.

Gwyn Williams prepared a revised pre-proposal for the joint VCU/Jefferson Lab collaboration concerned with the re-commissioning of the Helios synchrotron for nanofabrication studies.  This was submitted to the Commonwealth Technology Research Fund office on June 14th.

A “surprise” safety inspection of the FEL facility was held June 22 by representatives of the DOE Site Office, DOE’s Area Office in Oak Ridge and OSHA.  No violations were found and good marks were given for our training and procedures.  A useful recommendation was made for following up on the disposition of chemicals and supplies that are brought to the FEL facility by users.

Project Cost Performance:
The project (Phase 1) budget for the period June 1, 2000 to September 30, 2001 is $9,029k.  The Phase 2 project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002.  The project through the month of June has a total of $4,419k of performance scheduled (assuming the project started at the originally planned start date of April 1, 2000).  The work performed through the end of June was $4,904k, which is 54% complete vs 49% scheduled.  The actual cost accrued through June totals $6,130k.  This results in a schedule variance of +$485k and a cost variance of -$1,226k. We are on track for having the first phase of the project fully obligated by September.  (The total project cost to date-actual costs plus commitments is $7,782 k).  The negative cost variance presently showing is largely an artifact of having to obligate and cost funds early in the project in order to ensure full obligation and costing by the project end date.

WBS 3 (Beam Physics):
Design team members participated in the Navy high power FEL workshop.  It was entertaining, particularly when the "ship" guy said, essentially, "consider a cubical FEL".  1 MW looks, however, within the realm of reason, and, with proper coaxing, the upgrade (with a completely new injector and using appropriately configured 7-cell linac SRF) might someday run to near MW class.

Technical activities centered on analysis of ancient (Oct 1999) difference orbit data.  Reasonable agreement between model and measurement is being achieved.  This is unfortunate, because the difference orbit data results are anti-correlated with predictions & measurements of BBU thresholds.  Oops. More to follow.

The talk given at the 2001 Particle Accelerator Conference on the Upgrade design can be found at :  http://www.jlab.org/~douglas/FELupgrade/talks/PAC2001.ppt

We are exploring why the dumped beam is lower in energy than the injected energy (it’s because we power recover, not energy recover).  In the process, it became apparent that the power (not energy) recovery process utilizes only about half the available longitudinal aperture in the linac; a rebaseline of the process to energy recover the central energy of the distribution will require use of additional RF forward power, but
should allow recovery of much larger momentum spread at the dump without loss.

WBS 4 (Injector):
Kovar rings for the ceramic stack assembly have been welded into flanges and the mid-voltage ring, and the assembly was leak checked. All of the components for the ceramic stack went through final cleaning before the stack is welded together.

The shop is ordering pipe and tubing for the gun chamber.  The gun stand and a handling device for the ceramic stack are being designed.

Worked on the FET system to obtain preliminary results for nitrogen implanted titanium.

The upgraded HVPS unit was received for the gun. It will be installed next year.  A new Gun Tank and Transmission Line elbow are being designed now.  Continued refining the gun chamber drawings which should be complete with the exception of the screens and associated mounts which are 50% complete.  The materials for the gun chamber are in house.

WBS 5 (SRF):
Space frame and vacuum vessel punch list generated with vendor.  Planned delivery for both 1 July.
The cavity team tested a 7 cell cavity structure which had previously shown a degraded Q symptomatic of Q disease.  Following a bakeout, the follow-up performance test showed improved Qs.  There is some concern that the niobium used for the recent cavity tests may be more susceptible to Q disease because it has a noticeably smaller grain structure than previously used niobium.  A careful comparison of the chemical and physical characteristics of our existing niobium stocks is being made.

Received prototype vacuum vessel and spaceframe.  Acceptance inspection in progress.  Received reworked spaceframe tooling.  Required repair/rework of aluminum welds.

WBS 6 (RF):
Zone 3 –
o The circulators for this zone were staged.  They will be installed within a few weeks.
o B. Nichols was assigned to complete this zone by the end of August.  The Circulators are staged for installation.  The HVPS is due to be received in 2 – 3 weeks from Hipotronics.

Zone 4 –
o The 2.5 watt amplifiers were replaced with new 5 watt versions.  The new amplifiers will be tested when   RF Control modules are available.  The missing circulator has been staged, but not yet installed.
o No progress on completing the punch list.  New RF Control Modules are being built, but none have been received to date.

Injector HVPS –
o RFQ's for the transformers and inductors to upgrade the HVPS for the Quarter were sent to 7 prospective vendors.
Injector Klystrons –
o CPI believes they have resolved the helium leak test questions are proceeding to complete the three 100 kW klystrons.  No date has been set for factory tests.
Quarter HVPS's –
o Quotes are due today for the transformers and inductors to upgrade the HVPS's.  Receiving information for the rectifier assemblies and SCR controllers.

WBS 8 (I&C):
Lots of LSS work.  The wiring of the hutches in labs 3 and 6 has been finished and all necessary limit switches and smoke detectors installed.  The Low Level certification of all labs is complete.  James and Dan assisted with the Alignment Mode certification as well. James set up a video system for the experiment in Lab 3. OBPM testing continues.  James is finishing the drawing of the daughter board to take to EECAD. Two significant additions were made to the website this week.  All FEL Drawing numbers can now be located on the web.  A search function allows searching by drawing number, title, or drawn by name.  All user lab phone numbers are also posted on the website.  Future uses may see experimenters listed by name as they are using the facility. Bids are in on the assembly of the Ion Pump Power Supplies.  They will be shipped out for assembly early next week.  The front panel for the new VME Timing Circuit has been ordered expected delivery date is early July.

OBPM boards were completed and a daughter sample and hold board designed and built for testing.  They are currently installed in the system and appear to be working as designed.  The "auto-lock" pico motors are being tested today. The beamviewer assembly area is almost ready to start production.  The clean area has been constructed and the electricians ran power to the ventilation hood and lights, so we expect this project to
begin soon.  Drawings are nearly complete for the LSS connections to the hutches in labs 3 and 6.  All labs have successfully completed high & low level certifications.  Several additional isolation and interface module VME cards have been built.  Some for testing the new timing circuit which is nearing completion, and some for spares to other circuits. An inventory of all available VME cards has been added, along with a search function to see the location and status of each card.  Media gallery is now part of a database making the
addition of new pictures much easier.  Anyone with questions on how to do this should contact Mike Aston.  Also in the works is a search program for the task system that will allow easy access to tasks that have already been completed.  This will provide a useful historical data base as well as a way to prevent "re-inventing the wheel" when two similar tasks are performed. An additional user/camera power supply was built. System development on the "Rabbit" board is progressing well. New drawings for the OBPM and the Absolute Value/Sample and Hold Board were given to EECAD this week.

The installed Optical Beam Position Monitors appear to be working and will provide a useful tool for steering and monitoring beam position. The two that were calibrated using the drive laser are installed in Can 6 (Optics Control) and Can 2(downstairs in the beam transport line).  More testing is planned as a more stable running mode is achieved. A User Power Supply was installed in Lab 1 for use by the folks working there.  It is providing a +5vdc trigger signal for their computer when the shutter leaves the shut position.  The computers they brought with them have been provided IP addresses so they can communicate with each other, but do not affect lab security, or the LAN.  An additional power supply is currently under construction.  Video was re-connected to the conference room television so the observers can watch from there as the control room was getting rather crowded. The servo motor box has been worked, but is not yet complete.  Color plots of the VME Timing/Delay Circuit board have been completed and are awaiting
review.  Components are on order to support assembly once the board is manufactured.  The assembly contract for the Ion Pump Power Supplies was awarded to Assembly and Design, Inc. in Richmond and parts were delivered Thursday.  We hope to see the early results back soon. The assembly drawing for the controller board of the Ion Pump Power Supply was signed and submitted to document control prior to parts being sent to Assembly and Design, Inc.  Also purchased an audio system to allow audio recording of experiments as well as video.  New VME Timing/Delay boards are on order.  Mike and Patrick have been working on the computer in User Lab 4 to get it connected to the user LAN.  Mike has added a search
feature to the Task System on the laser website.  The next step in this process will be the ability to search the FLOG in a read only format.  Patrick has programmed the rabbit board to operate servo motors remotely and built a chassis for it.  The next step will be to access the LAN for control from off site.

The pulse truncater for the Drive Laser has been connected.  James and Dan have completed the video patch panel for the hutch in User Lab 3.  The spare VME Interface modules are complete.  The Isolation modules are waiting on parts.  OBPM work continues.  A tracking system using serial numbers has been established for better control of testing of the detectors.  The absolute value/sample and hold board drawing is in EECAD.  Prototype boards should be out soon.

Drawings in EECAD:
- Molectron Quad Buffer Amplifier, Fabrication, Schematic and Assembly;
- Analog/Differential Driver Board Schematic;
- Pyro-electric Detector Buffer Amplifier, Schematic, Fabrication and Assembly;
- Ion Pump Power Supply Controller Board Schematic;
- Beam Viewer Chassis Test Box Schematic, Assembly and Fabrication;
- LSS PLC Test Box Schematic, Assembly and Fabrication;
- User Shutter Control Box Wiring Diagram;
- MPS System Diagram;
- Absolute Value/Sample and Hold Board Schematic, Assembly and Fabrication;
- LSS Wiring Diagram- Revision.

WBS 9 (Transport):
Injector Dipoles (DU/DV):
o The drawings of both magnets were checked and sent back to DULY research for revisions.
Optical Chicane Dipole (DW):
o We started working on the specifications for magnet manufacturing and assembly.
Arc Dipoles (GY, GX, GQ):
o Advanced Energy Systems worked on and finished the details of the leads, bus bars, manifold, alignment cartridge mounts and lead cover of the Arc Bend Dipole (GX).  All these parts have to be able to work in both a left hand configuration as well as a right hand for two positions and the left hand configuration when the magnet is flipped over for the remaining two positions.   They submitted all drawings of this magnet for checking.
o Comparisons between the symmetric Preliminary GX and the present axisymetric design, showed that the central flat field zone shifts toward the narrow end by a few cm but remains nearly identical at about 57 % of the steel length at the highest field, thus cross checking the TOSCA and ANSYS.
o The IR Demo is successful at about half field where the +-1 part in 10,000 flatness extends to 65% of the length.   AES is running a magnetic model for the upgrade at 145 MeV/c (near the field the machine will run for its commissioning and early running).  We define the operating point at this field and evaluate all our results on all the magnets at this momentum.  I similar process was used on the IR Demo
o In magnetic modeling and pursuit of the flat longitudinal field above, the use of a Purcell gap or the widening of the field clamp gap failed.  Further work using the use of micrometric decreases in gap size in bands near the edges succeeded and we will be going on to now establish the effective length by adjusting the position of the end faces.  This will enable us to then determine the final gap and Purcell gap on the 180 Degree Dipole (GY).
o AES started design of the last magnet of the series, the Reverse Bend (GQ).  It will be able to feed directly off the design of the Regular Bend (GX), so rapid progress is expected.
3 inch quad (QX):
o The coil manufacturing contract was placed.
o The core manufacturing bids were received and evaluated.  A series of questions for the vendors to clarify their submissions were generated and responses received.  Final evaluation is in process
o The first multifaceted cuts on the chamfers of the ends of the poles reduced n =10 pole but the N=6 pole rose.  We are evaluating the results
Trim Quad (QX):
o Final magnetic modeling reduced the width by 2 inches.  A last pass at the drawings incorporated the changes.
Corrector Dipoles (DB & DJ):
o Manufacturing started.
Sextupole (SF):
o Robin Wines continued work on the 3D model in TOSCA.  Drafting design started.
Octupole (OT):
o Dave Douglas reduced the field required of the Octupole considerably.  A task order for DULY Research to start magnetic modeling and design was composed.
Beam Line and Vacuum:
o Substantial work was done on the Interface Control Document for the task definition for the Arc and Optical Chicane chambers.
o The layout drawings for the girders between cryomodules and for the return leg were completed and submitted to all interested parties for review.
o Butch Dillon Townes was able to advance the vacuum design by tabulating the vacuum pump needs of the Upgrade and publishing them.  A few areas need further evaluation but we now have numbers we can work with.
o The Dipole Steel order is still awaiting vendor response to a revised flatness tolerance that is due today.
o The copper order for all the major dipoles was placed

WBS 10 (Wiggler):
The tubing for the vacuum chamber has been received and the machine shop made up test pieces to test the welding technique.  Mike Beck has been assigned to manage the fabrication of the chambers.

Four end coils are now wound for the wigglers.  A meeting was held to finalize plans for the measurement of the wiggler. The 3D Hall probe was received.  A plan for calibrating the probe was accepted at the meeting.

Analysis of the dispersion section measurements with a single turn correction coil verify that the field integral can be zeroed to +/- 10 G-cm over the full range of the magnet strength.  An alternate scheme will be tested as well.  A scheme for adding a correction coil to the dispersion section was developed.  The correction coils for the wigglers are done and will be mounted next week.  A test weld piece was made of a viewer cross for the wiggler chamber.  Despite using a fusion weld technique, the measured permeability was approximately 1.5 at the weld bead.  Since this is much higher than for the IR Demo we decided to measure the effect of the weld on the dispersion section field.  Measurements were made which indicated changes to the field of about 1 G near the weld. Since the electron beam does not get very near the weld and since a change of 1 G over a few centimeters would be acceptable anyway, we decided that the weld technique was acceptable.  The full chambers will now be fabricated.  Most of the rest of the pieces for the viewer camera assemblies and the vacuum chamber piece parts are now fabricated and in house.

WBS 11 (Optics):
We have begun setting up a test stand for the mirror metrology system.  We are receiving components for the ultraviewers, part of the alignment system.

The optical cavity assembly design has progressed in the following areas: we've made changes to the placement of the relative cavity length lvdt, and added an attachment to the top of the gimbal yoke to facilitate installation an removal.  We've increased the size of the openings across the gimbal support frame to facilitate removal of the deformable mirrors.  K. Jordan kindly offered to supply a stepper motor that will be used for the y motion of the assembly.  Work is transitioning to the detailing phase, with calculation of loading on the rails (used for y and z motion), thickness of support structures, etc.  We received our Veeco laser interferometer, and scheduled installation for July 12 & 13.  We defined the assembly of parts for the outgassing tests. (Thanks to T. Siggins & D. Bullard)

With regard to the Upgrade:
o detailing began on the internal mounting hardware in the optical cavity chambers
o we received the second deformable mirror mount from the vendor; next step is to begin qualification tests using the Lab 2 laser
o we reviewed drawings for the 2 inch deformable mirror mount and sent them on for continuing detailing
o we began design of the insertable mirror for the diagnostics mounted between the outcoupler chamber and the collimator.  In the process, we began exploring a suggestion that we use two insertable viewers (or ultraviewers), a design which will result in cost savings.

We completed the low level certification of the LSS for all the user labs and the OCR, also completed the high level (final) certification for Labs 1 and 3 and the OCR.

We assisted the JTO-funded research team from NRL with initial checkout and integration of their diagnostics with our set-up in preparation for this month’s start of the JTO run for “lethality” measurements.  We installed the majority of our diagnostics.  We ordered more material for the experiments, and made arrangements to receive other materials from NRL.

The nanotube team (CWM, NASA, Penn State) was given approximately 20 hours of run time, which they used very successfully for nanotube production.  They believe that they have generated sufficient material for extensive purification and characterization of the fibers.

Two weeks of FEL run time were devoted to the JTO lethality measurements. By 17:00 on the last day of the first run week, we completed 15 sample suites (with 6 samples per suite).  Sample effects are being measured on metal, ceramic and composite samples at irradiances between 1-10 kW/cm2. Preliminary analysis shows sample melting on the stainless steel samples with the FEL beam where we would expect no melting with a cw laser beam at the same power and wavelength.  At the end of the second week of the JTO run more than 40 sample suites were run. Substantial amounts of data have been acquired and will require some significant time to fully reduce and analyze. On one day we ran for 1 kW for over 6 hours. We have also seen FEL efficiency run as high as >2% due to good optimization of the beam transport. This bodes well for our Upgrade plans which only require 0.8%.  Unfortunately with such a high CW power running uninterrupted the un-cooled mirrors in the optical transport line eventually heated to unacceptable levels and oxidized a reflective coating on one. Replacement of that mirror brought us back to full power operation the next morning.  As a consequence of that experience we are limiting our high power CW
running to periods of an hour or two until we can attach some temperature stabilization clamps to the most stressed components.  The machine ran well during all this activity and we are able to produce 3 more runs of nanotube production as well as parasitically map the Thompson X-ray production while the JTO data was being acquired. The fluxes produced there were also high enough to saturate that detector under many conditions.  We are adding tighter collimation.

Privacy and Security Notice